# Liming Liu1    Zhenghao Xu1    Zixuan Zhang1    Hao Kang1 Zichong Li1    Chen Liang2    Weizhu Chen2    Tuo Zhao1

### v2

@@ 3 COSMOS: A Hybrid Adaptive Optimizer
4:     $M_{t}\leftarrow\beta_{1}M_{t-1}+(1-\beta_{1})G_{t}$
5:     $B_{t}\leftarrow\texttt{NS5}(M_{t}/\|M_{t}\|_{\rm F})$
6:     $W_{t+1}\leftarrow W_{t}-\eta B_{t}\cdot\sqrt{m}$
7:  end for
Algorithm 3 MUON for an $m\times n$ layer $W$. Per layer, we maintain one matrix: $M\in\mathbb{R}^{m\times n}$.
Memory Usage Comparison: For comparison, we list the memory usage of the optimization states in Adam, Adam-mini, SOAP, MUON and COSMOS for training transformer models in Table 1. For simplicity, we assume that the attention weight matrices $W_{Q},W_{K},W_{V},W_{O}\in\mathbb{R}^{d\times d}$ and the MLP weight matrices $W_{1}\in\mathbb{R}^{d\times 4d}$ and $W_{2}\in\mathbb{R}^{4d\times d}$. Note that in practical LLMs, the dimensionalities of $W_{1}$ and $W_{2}$ might slightly vary. Moreover, we assume that the rank of the projection is $r=0.05d$ for COSMOS.
Table 1: Memory usage of the optimization states in different algorithms for training transformer models.
@@
Two-sided COSMOS. For simplicity, we only consider the one-side version of COSMOS in this paper. Like SOAP, COSMOS can be further generalized to a two-sided version. See more details in Appendix B.
@@ 4 Experiments
Models and Datasets. We train LLaMA-type models (Touvron et al., 2023) on the C4 dataset (Raffel et al., 2020), which is a colossal, cleaned version of Common Crawl’s web crawl corpus for pre-taining. We conduct comprehensive experiments and ablation studies on 130M models and demonstrate the token efficiency of COSMOS. We then scale up to 350M and 1B models to showcase the memory efficiency and small computational overhead of COSMOS. Due to limited computational resources, experiments on these larger models are less comprehensive, while still capable of illustrating the efficacy of our method. We train for one epoch on a portion of the C4 dataset, ranging from 5B to 26B tokens, and scaling with the model size according to the scaling law (Kaplan et al., 2020). We set the maximum sequence length as 1024 by default. Besides LLaMA models and C4 dataset, we also experiment with GPT-2 (Radford et al., 2019) on WikiText-103 dataset (Merity et al., 2016), presented in Section C.2.
@@
4.1 Comparison on LLaMA-130M
For LLaMA-130M, we train on a 5B subset of the C4 dataset. We compare COSMOS’s validation loss with that of Adam, Adam-mini, GaLore, MUON, and SOAP. Hyperparameters. For each method, we tune the corresponding learning rate to obtain optimal performance. We select the rank $r=64$ for COSMOS and $r=256$ for GaLore. To avoid multiple hyperparameter tuning, we set the discount factor in COSMOS as $\gamma=\eta/\eta_{0}$, where $\eta_{0}$ is the learning rate of Adam for training the embedding and output weights in the implementation of COSMOS. Other hyperparameter choices follow Zhao et al. (2024a) and are provided in detail in Section A.1.1.
@@ 4.1 Comparison on LLaMA-130M
Ablation on learning rates. We experiment with different learning rates while keeping the rank $r=64$ and discount factor $\gamma=\eta/\eta_{0}$ for COSMOS. As shown in the Table 3, COSMOS is not very sensitive to the learning rate, and it achieves the best performance at 5e-4. As a comparison, MUON is more sensitive to the learning rate, and it underperforms COSMOS across all learning rates.
@@
Table 4: Validation perplexity of COSMOS under different ranks $r$ and discount factors $\gamma$. Our method is not very sensitive to $r$ and $\gamma$, and consistently outperforms MUON (20.69) except for only one configuration ($r=128$, $\gamma=1$).
$r$\$\gamma$ 0.1 0.25 0.5 1
@@
Effect of normalization. COSMOS applies a normalization step (Line 9 in Algorithm 1) after the NS transformation. To illustrate that normalization alone cannot fully explain the COSMOS’s efficiency, we apply the same normalization to MUON and compare their performances. As shown in Figure 3, MUON with normalization performs similarly to the version without normalization, and both are worse than COSMOS. This implies that normalization does not contribute as much as other components of COSMOS, including the SOAP-like update in the leading eigensubspace. We present the same experiment with LLaMA-350M in Section C.1.
Figure 3: Comparison of COSMOS, MUON and MUON with normalization for LLaMA-130M on C4. MUON’s performance does not differ by much with normalization, implying that normalization is not the dominant reason why COSMOS outperforms MUON.
GaLore degradation on long sequences. In our experiment, we observe that GaLore performs much worse than COSMOS and other baselines including Adam. To validate the correctness of our implementation, we compare GaLore with COSMOS and Adam on 256 sequence length as adopted in Zhao et al. (2024a). As shown in Figure 2, GaLore and Adam are comparable in this shorter sequence setting, suggesting that GaLore degenerates on long sequences. This observation aligns with Liang et al. (2024), while a similar phenomenon appearing in fine-tuning is reported by Pan et al. (2024). In contrast, COSMOS consistently outperforms Adam from short to long sequences without suffering from degradation.
@@ 4.2 Scaling up to LLaMA-350M and LLaMA-1B
To further illustrate the efficacy of COSMOS, we scale up to larger models and more tokens. For LLaMA-350M, we train on a 10B subset of the C4 dataset for 5000 steps. We compare COSMOS with Adam, GaLore, MUON, and SOAP. For LLaMA-1B, we train on a 26B subset for 13000 steps. Given limited computation resources, we compare COSMOS with Adam and MUON. Adam serves as the standard baseline, while MUON achieves better performance than Adam while requiring less memory and little computational overhead. Although SOAP demonstrated superior performance among baselines, we exclude it from our comparison as its complete training for the 1B model exceeds our available resources. More experiment details are provided in Sections A.1.2 and A.1.3.
Main results of token efficiency. Figures 4 and 5 display the validation loss curves for the 350M and 1B models, and Table 2 presents their final validation perplexities. COSMOS demonstrates superior performance compared to all baselines across both model sizes, matching the results on the 130M model and showcasing consistent token efficiency across different model sizes.
Figure 4: Performance of LLaMA-350M trained on C4 dataset. COSMOS consistently outperforms baseline methods.
Figure 5: Performance of LLaMA-1B trained on C4 dataset. COSMOS consistently outperforms Adam and MUON.
Memory and computation time profiling. To illustrate the memory efficiency and small computation overhead of COSMOS, we conduct a profiling experiment on LLaMA-1B to track memory usage and wall-clock time. We fix the batch size 10 and gradient accumulation step 25 for all methods for a fair comparison. We record the maximum GPU memory usage and time spent during the entire forward-backward propagation and optimizer update process for one iteration. As shown in table 5, COSMOS achieves much lower maximum GPU memory usage than Adam (6.8%) and SOAP (19.4%), with slightly more overhead compared to MUON. In terms of wall-clock time per iteration, COSMOS is comparable to MUON and is much better than SOAP. The fastest Adam method cannot achieve the same level of token efficiency as COSMOS. Therefore, COSMOS strikes a good balance between token and memory/computation overheads, achieving the best final perplexity at a much lower cost.
Table 5: GPU memory usage and wall-clock time per iteration on 1B model. We fix the batch size to be $10$ for all methods. COSMOS has significantly less memory usage than SOAP and is comparable to memory-efficient methods like Adam-mini and MUON, without introducing much computation overhead.
@@
Table 6: System performance on single NVIDIA A100-80G GPU and corresponding throughput (number of samples processed per second on C4 dataset) of 1B model. Max batch size is defined as the maximum number of samples that fit within the GPU’s memory capacity. Throughput is reported as the number of samples the GPU processes per second (samples/s)
@@
Throughput. To better compare the methods in practical setting, we evaluate the maximal batch size and throughput of COSMOS and baselines. We experiment on a single NVIDIA A100 GPU with 80GB memory. The input sequence length is set to 1024. As shown in Table 6, COSMOS is 10.8% faster than SOAP and comparable to MUON.
4.3 Summary of Numerical Experiments
We summarize our main experiment results. (1) COSMOS achieves the best token efficiency comparable to SOAP. (2) Our memory usage and throughput are significantly better than SOAP. (3) Our method performs consistently across different model sizes, different sequence lengths, and different hyperparameter configurations.
@@ 5 Conclusion
Appendix A Experiment Details
@@
[…]
@@ A.1.1 Tuning on 130M model
[…]
For GaLore, We follow the setting in Zhao et al. (2024a), set rank=256, and scale factor $\alpha=0.25$. The projection update frequency is 200 for 20k training steps, thus we decrease it to 50 for our 5k training steps. According to Zhao et al. (2024a), the learning rate should be larger than Adam’s.
[…]
@@ A.1.2 Tuning on 350M model
[…]
@@ A.1.3 Tuning on 1B Model
[…]

### v3

@@ 3 COSMOS: A Hybrid Adaptive Optimizer
4:  $M_{t}\leftarrow\mu M_{t-1}+G_{t}$
5:  $N_{t}\leftarrow\mu M_{t}+G_{t}$
6:  $B_{t}\leftarrow\texttt{NS5}(N_{t}/\|N_{t}\|_{\rm F})$
7:  $W_{t+1}\leftarrow W_{t}-\eta B_{t}\cdot\sqrt{m}$
8:end for
Algorithm 3 MUON
Remark 1.
As can be seen, COSMOS only needs to maintain four matrices in the memory: $M_{t}\in\mathbb{R}^{m\times n}$, $U_{t}\in\mathbb{R}^{n\times r}$, $S_{t}\in\mathbb{R}^{r\times r}$ and $V_{t}\in\mathbb{R}^{m\times r}$. In sharp contrast, even one-sided SOAP (2) needs to maintain $M_{t}\in\mathbb{R}^{m\times n}$, $L_{t}\in\mathbb{R}^{n\times n}$, $U_{t}\in\mathbb{R}^{n\times n}$ and $V_{t}\in\mathbb{R}^{m\times n}$. The resulting memory overhead is which is significantly larger than that of COSMOS.
Remark 2.
Recall that the the computation complexity of the QR decomposition on a matrix of the shape $n\times r$ is $O(nr^{2})$ when $rn$, so the low rank QR decomposition of $\beta_{2}U_{t-1}S_{t-1}+(1-\beta_{2})G_{t}^{\top}G_{t}U_{t-1}$ in COSMOS is actually very quick since $r\ll n$ (and much quicker than that in SOAP, which is $O(n^{3})$). Therefore, unlike SOAP which needs to consider the preconditioning frequency for performing QR decomposition, we can carry out QR decomposition at every step with virtually no overhead. In addition, PyTorch provides an efficient implementation of QR method, which is also used by SOAP. In Table 5, we provide the comparison of wall-clock time per iteration to show that compared to MUON, COSMOS only incurs a very slight increase in wall-clock time.
3.1 Memory Usage Comparison
For comparison, we list the memory usage of the optimization states in Adam, Adam-mini, SOAP, MUON and COSMOS for training transformer models in Table 1. For simplicity, we assume that the attention weight matrices $W_{Q},W_{K},W_{V},W_{O}\in\mathbb{R}^{d\times d}$ and the MLP weight matrices $W_{1}\in\mathbb{R}^{d\times 4d}$ and $W_{2}\in\mathbb{R}^{4d\times d}$. Note that in practical LLMs, the dimensionalities of $W_{1}$ and $W_{2}$ might slightly vary. Moreover, we assume that the rank of the projection is $r=0.05d$ for COSMOS.
Table 1: Memory usage of the optimization states in different algorithms for training transformers.
@@ 4 Experiments
Models and datasets. We train LLaMA-type models (Touvron et al., 2023) on the C4 dataset (Raffel et al., 2020), which is a colossal, cleaned version of Common Crawl’s web crawl corpus for pre-taining. We conduct comprehensive experiments and ablation studies on 130M models and demonstrate the token efficiency of COSMOS. We then scale up to 350M and 1B models to showcase the memory efficiency and small computational overhead of COSMOS. Due to limited computational resources, experiments on these larger models are less comprehensive, while still capable of illustrating the efficacy of our method. We train for one epoch on a portion of the C4 dataset, ranging from 5B to 26B tokens, and scaling with the model size according to the scaling law (Kaplan et al., 2020). We set the maximum sequence length as 1024 by default.
Besides LLaMA models and C4 dataset, we also conducted experiments with modded-NanoGPT (Jordan et al., 2024) on FineWeb (Penedo et al., 2024) and GPT-2 (Radford et al., 2019) on WikiText-103 (Merity et al., 2016) to evaluate the effectiveness of COSMOS across different settings.
4.1 Comparison on LLaMA-130M
For LLaMA-130M, we train on a 5B subset of the C4 dataset. We compare COSMOS’s validation loss with that of Adam, Adam-mini, GaLore, MUON, and SOAP.
Hyperparameters. For each method, we tune the corresponding learning rate to obtain optimal performance. We select the rank $r=64$ for COSMOS and $r=256$ for GaLore. To avoid multiple hyperparameter tuning, we set the discount factor in COSMOS as $\gamma=\eta/\eta_{0}$, where $\eta_{0}$ is the learning rate of Adam for training the embedding and output weights in the implementation of COSMOS. Other hyperparameter choices follow Zhao et al. (2024a) and are provided in detail in Section A.1.1.
@@ 4.1 Comparison on LLaMA-130M
Ablation on learning rates. We experiment with different learning rates while keeping the rank $r=64$ and discount factor $\gamma=\eta/\eta_{0}$ for COSMOS. As shown in the Table 3, COSMOS is not very sensitive to the learning rate, and it achieves the best performance at 5e-4. As a comparison, MUON is more sensitive to the learning rate, and it underperforms COSMOS across all the learning rates.
@@
Table 4: Valid perplexity of COSMOS under different $r$ and $\gamma$. COSMOS is not very sensitive to $r$ and $\gamma$, and consistently outperforms MUON (20.69) except for only one config ($r=128$, $\gamma=1$).
[…]
@@
Effect of normalization. COSMOS applies a normalization step (Line 9 in Algorithm 1) after the NS transformation compared to MUON (Algorithm 3). Empirically, we find that COSMOS also outperforms the normalized version of MUON (see Figure 5 in Section D.1). This implies that normalization is not the only driving force behind COSMOS’s efficiency.
GaLore degradation on long sequences. In our experiment, we observe that GaLore performs much worse than COSMOS and other baselines, including Adam. Such a degradation is less significant on the shorter sequences with length 256 (see Figure 8 in Section D.5, the setting adopted in the original GaLore paper (Zhao et al., 2024a). This observation of GaLore degradation on long sequences aligns with Liang et al. (2024), while a similar phenomenon appearing in fine-tuning is reported by Pan et al. (2024). In contrast, COSMOS consistently outperforms Adam from short to long sequences without suffering from degradation.
@@ 4.2 Scaling Up to LLaMA-350M and LLaMA-1B
To further illustrate the efficiency of COSMOS, we scale up to larger models and more tokens. For LLaMA-350M, we train on a 10B subset of the C4 dataset for 5000 steps. We compare COSMOS with Adam, Adam-mini, GaLore, MUON, and SOAP optimizers. For LLaMA-1B, we train on a 26B subset for 13000 steps. Given limited GPU resources, we compare COSMOS with Adam and MUON. Adam serves as the standard baseline, while MUON achieves better performance than Adam while requiring less memory and little computational overhead. Although SOAP show superior performance among baselines, we exclude it from our comparison as its complete training for the 1B model exceeds our available resources. More experiment details are provided in Sections A.1.2 and A.1.3.
Main results of token efficiency. Figures 2(a) and 2(b) display the validation loss curves during training for the 350M and 1B models, and Table 2 presents their final validation perplexities. COSMOS demonstrates superior performance compared to all baselines across both model sizes, matching the results on the 130M model and showcasing consistent token efficiency across different model sizes.
(a) LLaMA-350M trained on C4 dataset.
(b) LLaMA-1B trained on C4 dataset.
Figure 2: Comparison of performance on LLaMA-350M and LLaMA-1B trained on the C4 dataset.
Memory and computation time profiling. To illustrate the memory efficiency and small computation overhead of COSMOS, we conduct a profiling experiment on the 1B model. We fix the batch size as 10 and the gradient accumulation steps as 25 for all methods and record the maximum GPU memory usage and time spent during the entire forward-backward propagation and optimizer update process for one iteration. As shown in Table 5, COSMOS achieves much lower maximum GPU memory usage than Adam (6.8%) and SOAP (19.4%), with slightly more overhead compared to MUON. In terms of wall-clock time per iteration, COSMOS is comparable to MUON and is much better than SOAP. The fastest Adam method cannot achieve the same level of token efficiency as COSMOS. Therefore, COSMOS strikes a good balance between token and memory/computation overheads, achieving the best final perplexity at a much lower cost.
To better compare the methods in a practical setting, we evaluate the maximal batch size and throughput of COSMOS and baselines on a single NVIDIA A100 GPU with 80GB memory. The input sequence length is set to 1024. As shown in Table 6, COSMOS is 10.8% faster than SOAP and comparable to MUON.
Table 5: GPU memory usage and wall-clock time per iteration on 1B model. We fix the batch size to be $10$ for all methods. COSMOS has significantly less memory usage than SOAP and is comparable to memory-efficient methods like MUON, without introducing much computation overhead.
@@
Table 6: System performance on single NVIDIA A100-80G GPU and corresponding throughput (number of samples processed per second on C4 dataset) of 1B model. Max batch size is defined as the maximum number of samples that fit within the GPU’s memory capacity. Throughput is reported as the number of samples the GPU processes per second (samples/s).
@@
Wall-Clock time plot for LLaMA-1B. Based on the throughput we calculate in Table 6, we rescale the X-axis of Figure 2(b) to be wall-clock time and present the result in Figure 7 in Section D.4. Our results indicate that, in terms of actual training time, COSMOS still outperforms MUON and the Adam baseline, demonstrating COSMOS’s potential for efficient pretraining.
4.3 Additional Experiments on Other Settings
Most current works on pretraining optimizers, due to limitations in resources and time, focus on a single model architecture and a single dataset — for example, GaLore (Zhao et al. (2024a),LLaMA on C4), SOAP (Vyas et al. (2024), OLMo (Groeneveld et al., 2024) on C4) and Muon (Jordan et al. (2024), Modded-NanoGPT on FineWeb). Our experiments on LLaMA with C4 are already aligned with these prior works, validating the performance of COSMOS. To demonstrate that COSMOS retains its advantages under other settings as well, however, we also conduct experiments in the following additional settings:
Modded-NanoGPT on FineWeb: To further verify the advantage of COSMOS over Muon, we conduct experiments in Muon’s original setting, namely Modified-NanoGPT on FineWeb. Since Muon has already performed extensive hyperparameter tuning in this setting, we do not tune Muon again but use the provided reproducible log. For COSMOS, we simply align the learning rate with that of Muon and follow other settings. We present the results in Figure 3. See Appendix B for the detailed configuration.
(a) 124M Modded-NanoGPT model trained on FineWeb dataset.
(b) 350M Modded-NanoGPT model trained on FineWeb dataset.
Figure 3: Comparison of optimization performance on 124M and 350M Modded-NanoGPT trained on the FineWeb dataset.COSMOS consistently outperforms MUON.
GPT2 on Wikitext-103: We also trained GPT2-small and GPT2-medium on WikiText-103 to compare COSMOS with Adam and Muon. In this setting, COSMOS still outperforms Muon and Adam. We present the result in Figure 4. See Section D.2 for the detailed configuration.
(a) GPT2-small trained on WikiText-103.
(b) GPT2-medium trained on WikiText-103.
Figure 4: Comparison of COSMOS, MUON and Adam on WikiText-103 using GPT2-small and GPT2-medium models. COSMOS consistently outperforms MUON and Adam.
@@ 5 Conclusion
Appendix A Experiment Details on LLaMA Models
@@
Default hyperparameters. In all algorithms, we choose first order momentum $\beta_{1}=0.9$ to align with and get a fair comparison with Adam baseline. We choose second order momentum $\beta_{2}=0.98$, which is also a widely used configuration after Liu et al. (2019) mentioned that it provides better training stability than 0.999. We set smoothing term $\epsilon=\text{1e-8}$ to align with the standard hyperparameter choice. We use the linear learning rate schedule to decay the learning rate to 0. To align with Zhao et al. (2024a), we set the warmup ratio to be 10% and weight decay to be 0.
@@ A.1.1 Tuning on 130M model
For Adam, we tune the learning rate on $\{$2.5e-4, 5e-4, 1e-3, 2e-3, 4e-3, 8e-3}\}. In our experiments, 2e-3 is the optimal learning rate and 8e-3 diverges. Then for SOAP, we also tune the learning rate on $\{$5e-4, 1e-3, 2e-3, 4e-3}\}. For Adam-mini, we just use the optimal learning rate of Adam, which is also 2e-3.
For GaLore, We follow the setting in Zhao et al. (2024a), set rank=256, and scale factor $\alpha=0.25$. According to Zhao et al. (2024a), the learning rate of Galore should be larger than Adam’s. They mentioned that Galore is not sensitive to hyperparameter and they use the same learning rate 1e-2 for all size of models after tuning, we simply tune galore in a range near 1e-2, which is $\{$5e-3, 1e-2, 2e-2, 4e-2}\}. The projection update frequency is 200 for 20k training steps, thus we decrease it to 50 for our 5k training steps.
For the implementation of MUON and COSMOS, the embedding and output layer will use Adam while other parts will use MUON/COSMOS algorithm. To avoid multiple hyperparameter tuning, we fix the learning rate for embedding and output layer to 2e-3, which is the optimal learning rate for Adam, and only tune the learning rate of hidden layers, whose optimizer is MUON/COSMOS. To be more specific, we tune the learning rate of hidden layers on $\{$1e-4, 2e-4, 5e-4, 1e-3, 2e-3}\}. It is worth noting that (Liu et al., 2025) suggests the optimal learning rate for MUON should be 0.2–0.4 times that of the Adam learning rate used for the embedding layer (2e-3 in our setting), which exactly falls within the range we searched.
@@ A.1.2 Tuning on 350M model
For Adam and SOAP, we tune the learning rate on $\{$2.5e-4, 5e-4, 1e-3, 2e-3, 4e-3, 8e-3}\}, which is same as the range in Section A.1.1. For GaLore, we set the rank to be 384, projection update frequency to be 50, and scale factor $\alpha=0.25$. Then we tune the learning rate of GaLore on $\{$5e-3, 1e-2, 2e-2, 4e-2}\}, which is also same as what we do in Section A.1.1.
For the implementation of MUON and COSMOS, we still fix the learning rate for embedding and output layer to be 2e-3 and only tune the learning rate of MUON/COSMOS for hidden layers. For MUON and COSMOS, we tune the learning rate on $\{$1e-4, 2e-4, 5e-4, 1e-3, 2e-3}\} to align with our setting in Section A.1.1. Also for COSMOS, we still set $\gamma$ to be the ratio of the learning rate of hidden layers to the learning rate of the embedding layer (which is fixed at 2e-3).
@@ A.1.3 Tuning on 1B model
We do not have enough resources to tune hyperparameters carefully on the 1B model. For Adam, we first try learning rate $\eta=\text{2e-3}$, but an extremely large loss spike occurred in the early stage. Then we decrease $\eta$ to 1e-3 and get the baseline result. For MUON and COSMOS, we still fix the learning rate for embedding and output layer to be 2e-3 and tune their learning rate on $\{$2e-4, 5e-4}\}. For COSMOS, we still set $\gamma$ to be the ratio of the learning rate of hidden layers to the learning rate of the embedding layer.
A.1.4 Discussion on learning rate tuning
We are discussing the learning rate used by MUON in their reproducible logs here to demonstrate that our learning rate falls within a reasonable range.
There are two versions of Muon implemented in the reproducible logs of modded nanogpt. In the early versions, the algorithm they used was consistent with what we described in Algorithm 3. This algorithm has been used on both 124M and 1.5B GPT models and achieved SOTA performance.
In this version, they used Adam’s baseline learning (3.6e-3) rate as the learning rate for the embedding and output layers on a 124M model, and used 3.6e-4 as the cleaning rate for Muon. In our experiment, since the Adam baseline learning rate we obtained was 2e-3, which is a little different with 3.6e-3, we also use this learning rate as the learning rate for the embedding and output layers. We avoid adjusting the learning rates of the embedding and output layers, as this would result in the tuning of both learning rates for two parts of parameters. Generally speaking, this would yield better results than adjusting only one learning rate, but this effect is not fair compared to the Adam algorithm with only one learning rate. For Muon’s learning rate in our experiments, our traversal set $\{$1e-4, 2e-4, 5e-4, 1e-3, 2e-3}\} is also relatively close to their 3.6e-4.
In the later reproducible logs, they made a simple modification to Muon, but did not mention whether this would improve the effect. This modification is to change the line 7 in Algorithm 3 to be $W_{t+1}\leftarrow W_{t}-\eta B_{t}\cdot\sqrt{\frac{m}{n}}$.
Considering that $n$ for different matrix parameters in the same LLM are the same (e.g., 768 in the 124M GPT), this method is just a simple rescale. However, due to the current scale being reduced by $\sqrt{n}$ times, this method generally requires a larger learning rate. For example, in their subsequent logs, they used 0.02 This learning rate is nearly equivalent to using $0.02/\sqrt{768}\approx 7.2\mathrm{e}-4$ in the first version, which is also close to our traversal set $\{$1e-4, 2e-4, 5e-4, 1e-3, 2e-3}\}.
Since MUON did not specify which version would be more advantageous, we used the first version in our experiments, as it provided more reproducible logs. At the time we began experimenting with MUON, the second version was not yet available. For consistency, we therefore continued with the first version, which was also adopted in the subsequent work on MUON scaling (Liu et al., 2025).